# Explicit Second-Order Min-Max Optimization Methods with Optimal Convergence Guarantee

### v4

@@ 1.2 Notation and Organization
We use bold lower-case letters to denote vectors, as in $\mathbf{x},\mathbf{y},\mathbf{z}$. For a function $f(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}$, we let $\nabla f(\mathbf{z})$ denote the gradient of $f$ at $\mathbf{z}$. For a function $f(\cdot,\cdot):\mathbb{R}^{m}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ of two variables, $\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y})$ or $\nabla_{\mathbf{y}}f(\mathbf{x},\mathbf{y})$ to denote the partial gradient of $f$ with respect to the first variable or the second variable at the point $(\mathbf{x},\mathbf{y})$. We use $\nabla f(\mathbf{x},\mathbf{y})$ to denote the gradient at $(\mathbf{x},\mathbf{y})$ where $\nabla f(\mathbf{x},\mathbf{y})=(\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y}),\nabla_{\mathbf{y}}f(\mathbf{x},\mathbf{y}))$ and $\nabla^{2}f(\mathbf{x},\mathbf{y})$ to denote the full Hessian at $(\mathbf{x},\mathbf{y})$. We write $\|\mathbf{x}\|$ for its $\ell_{2}$-norm. Finally, we use $O(\cdot),\Omega(\cdot)$ to hide absolute constants which do not depend on problem parameters, and $\tilde{O}(\cdot),\tilde{\Omega}(\cdot)$ to hide absolute constants and additional log factors.
The remainder of the paper is organized as follows. In Section 2, we present the setup of min-max optimization and provide the definitions and optimality criteria. In Section 3, we present an explicit yet conceptual second-order min-max optimization method and prove that it achieves an order-optimal iteration complexity of $\Theta(\epsilon^{-2/3})$. In Section 4, we propose a class of second-order min-max optimization methods with inexact second-order information and inexact subproblem solving, and we provide a crisp characterization of the complexity of solving each subproblem. In addition, we propose a set of subsampled Newton methods for solving the finite-sum min-max optimization problems. In Section 5, we present the results of experiments on real and synthetic datasets that demonstrate the practical efficiency of the proposed methods.
@@ 2 Preliminaries
A function $f$ is $\rho$-Hessian Lipschitz if $\|\nabla^{2}f(\mathbf{z})-\nabla^{2}f(\mathbf{z}^{\prime})\|\leq\rho\|\mathbf{z}-\mathbf{z}^{\prime}\|$ for all $\mathbf{z},\mathbf{z}^{\prime}\in\mathbb{R}^{n}$.
@@
$\begin{array}[]{rl}f(\mathbf{x}^{\prime},\mathbf{y})\geq f(\mathbf{x},\mathbf{y})+(\mathbf{x}^{\prime}-\mathbf{x})^{\top}\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y}),&\textnormal{for }\mathbf{x}^{\prime},\mathbf{x}\in\mathbb{R}^{m}\textnormal{ and any fixed }\mathbf{y}\in\mathbb{R}^{n},\\ f(\mathbf{x},\mathbf{y}^{\prime})\leq f(\mathbf{x},\mathbf{y})+(\mathbf{y}^{\prime}-\mathbf{y})^{\top}\nabla_{\mathbf{x}}f(\mathbf{x},\mathbf{y}),&\textnormal{for }\mathbf{y}^{\prime},\mathbf{y}\in\mathbb{R}^{n}\textnormal{ and any fixed }\mathbf{x}\in\mathbb{R}^{m}.\end{array}$
We also define the notion of global saddle points for the problem in Eq. (1.1).
@@
The existence of a global saddle point $\mathbf{z}^{\star}=(\mathbf{x}^{\star},\mathbf{y}^{\star})$ under Assumption 2.4 guarantees that $f(\mathbf{x}^{\star},\mathbf{y})\leq f(\mathbf{x}^{\star},\mathbf{y}^{\star})\leq f(\mathbf{x},\mathbf{y}^{\star})$ for all $\mathbf{x}\in\mathbb{R}^{m}$ and $\mathbf{y}\in\mathbb{R}^{n}$. Thus, we can adopt a restricted gap function (Nesterov, 2007) to provide a performance measure for the optimality of $\hat{\mathbf{z}}=(\hat{\mathbf{x}},\hat{\mathbf{y}})$ in the unconstrained convex-concave setting [^2] The restricted gap is also related to the classical Nikaidô-Isoda function (Nikaidô and Isoda, 1955) defined for a class of noncooperative convex games in a more general setting..
@@
$\mathrm{gap}(\hat{\mathbf{z}},\beta)=\max_{\mathbf{y}:\|\mathbf{y}-\mathbf{y}^{\star}\|\leq\beta}f(\hat{\mathbf{x}},\mathbf{y})-\min_{\mathbf{x}:\|\mathbf{x}-\mathbf{x}^{\star}\|\leq\beta}f(\mathbf{x},\hat{\mathbf{y}})$
where $\beta$ is sufficiently large such that $\|\hat{\mathbf{z}}-\mathbf{z}^{\star}\|\leq\beta$. Clearly, we have $\mathrm{gap}(\hat{\mathbf{z}},\beta)\geq 0$ since $f(\mathbf{x}^{\star},\hat{\mathbf{y}})\leq f(\hat{\mathbf{x}},\mathbf{y}^{\star})$ must hold true.
@@
A point $\hat{\mathbf{z}}=(\hat{\mathbf{x}},\hat{\mathbf{y}})$ is an $\epsilon$-global saddle point of a convex-concave function $f(\cdot,\cdot)$ if $\mathrm{gap}(\hat{\mathbf{z}},\beta)\leq\epsilon$. If $\epsilon=0$, it is a global saddle point.
@@
In our convergence analysis, we define the vector $\mathbf{z}=(\mathbf{x},\mathbf{y})\in\mathbb{R}^{m+n}$ and the operator $F:\mathbb{R}^{m+n}\mapsto\mathbb{R}^{m+n}$ as follows,
@@
In the following lemma, we provide the properties of the operator $F$ in Eq. (2.2) and its Jacobian $DF$ in Eq. (2.3) under Assumption 2.4 and 2.5. We note that most of the results in the following lemma are well known (Nemirovski, 2004) so we omit their proofs.
@@
Let $F(\cdot)$ and $DF(\cdot)$ be defined as above and Assumptions 2.4 and 2.5 hold true. Then, we have
@@
    $DF$ is $\rho$-Lipschitz continuous, i.e., $\|DF(\mathbf{z})-DF(\mathbf{z}^{\prime})\|\leq\rho\|\mathbf{z}-\mathbf{z}^{\prime}\|$.
@@
$(DF(\mathbf{z})-DF(\mathbf{z}^{\prime}))\mathbf{h}=\begin{bmatrix}I_{m}&\\ &-I_{n}\end{bmatrix}(\nabla^{2}f(\mathbf{z})-\nabla^{2}f(\mathbf{z}^{\prime}))\mathbf{h}.$ (2.4)
This implies that $\|(DF(\mathbf{z})-DF(\mathbf{z}^{\prime}))\mathbf{h}\|=\|(\nabla^{2}f(\mathbf{z})-\nabla^{2}f(\mathbf{z}^{\prime}))\mathbf{h}\|$. Thus, we have
$\|DF(\mathbf{z})-DF(\mathbf{z}^{\prime})\|=\sup_{\mathbf{h}\neq 0}\left\{\tfrac{\|(DF(\mathbf{z})-DF(\mathbf{z}^{\prime}))\mathbf{h}\|}{\|\mathbf{h}\|}\right\}=\sup_{\mathbf{h}\neq 0}\left\{\tfrac{\|(\nabla^{2}f(\mathbf{z})-\nabla^{2}f(\mathbf{z}^{\prime}))\mathbf{h}\|}{\|\mathbf{h}\|}\right\}.$
@@ 3 Conceptual Algorithm and Convergence Analysis
As a warm-up, we describe the scheme of Newton-MinMax which is a second-order version of the method in Lin and Jordan (2024) for min-max optimization and which yields an optimal global rate of $\Theta(\epsilon^{-2/3})$. We emphasize that Newton-MinMax is a conceptual algorithmic framework in the sense that it requires exact second-order information and requires the cubic regularized subproblem to be solved exactly.
@@ 3.1 Algorithmic scheme
Our method is a generalization of first-order extragradient method. Indeed, the $k^{\textnormal{th}}$ iteration consists of two important algorithmic components:
  • Gradient update: Compute $\Delta\mathbf{z}_{k}\in\mathbb{R}^{m}\times\mathbb{R}^{n}$ such that it is an exact solution of the nonlinear equation problem given by
    Compute $\mathbf{z}_{k+1}=\hat{\mathbf{z}}_{k}+\Delta\mathbf{z}_{k}$.
  • Extragradient update: Compute $\hat{\mathbf{z}}_{k+1}=\hat{\mathbf{z}}_{k}-\lambda_{k+1}F(\mathbf{z}_{k+1})$.
As suggested by Lin and Jordan (2024), we choose to update $\lambda_{k}$ in an adaptive manner and then prove that our method can achieve an order-optimal iteration complexity of $O(\epsilon^{-2/3})$ under Assumptions 2.4 and 2.5. Intuitively, such a strategy makes sense; indeed, $\lambda_{k}$ is the step size and would be better to increase as the iterate $\mathbf{z}_{k}$ approaches the set of global saddle points where the value of $\|\Delta\mathbf{z}_{k}\|$ measures the closeness. From a practical viewpoint, Algorithm 1 serves as an alternative to the current pipeline of line-search-based methods – which it simplifies by removing the need for an implicit binary search.
@@ 3.2 Convergence analysis
We provide our results on the iteration complexity of Algorithm 1 in the following theorem.
@@
Remark 3.2
Theorem 3.1 shows that Algorithm 1 has achieved the lower bound established in the literature for second-order VI methods (Lin and Jordan, 2024) and is thus order-optimal in this regard; in addition, it improves on the state-of-the-art bounds of Monteiro and Svaiter (2012); Bullins and Lai (2022); Jiang and Mokhtari (2022) by shaving off all logarithmic factors.
We define a Lyapunov function for Algorithm 1: $\mathcal{E}_{t}=\tfrac{1}{2}\|\hat{\mathbf{z}}_{t}-\mathbf{z}_{0}\|^{2}$ and use it to prove technical results that pertain to the convergence of Algorithm 1. The first lemma gives us a key descent inequality.
@@
$\sum_{k=1}^{t}\lambda_{k}(\mathbf{z}_{k}-\mathbf{z})^{\top}F(\mathbf{z}_{k})\leq\mathcal{E}_{0}-\mathcal{E}_{t}$
$+\underbrace{\sum_{k=1}^{t}\lambda_{k}(\mathbf{z}_{0}-\mathbf{z})^{\top}F(\mathbf{z}_{k})}_{\textbf{I}}+\underbrace{\sum_{k=1}^{t}\lambda_{k}(\mathbf{z}_{k}-\hat{\mathbf{z}}_{k})^{\top}F(\mathbf{z}_{k})-\tfrac{1}{2}\|\hat{\mathbf{z}}_{k}-\hat{\mathbf{z}}_{k-1}\|^{2}}_{\textbf{II}}.$
@@
II $\leq$ $\sum_{k=1}^{t}\left(\tfrac{13\lambda_{k}\rho}{2}\|\Delta\mathbf{z}_{k-1}\|^{2}\|\|\hat{\mathbf{z}}_{k-1}-\hat{\mathbf{z}}_{k}\|-\tfrac{1}{2}\|\hat{\mathbf{z}}_{k-1}-\hat{\mathbf{z}}_{k}\|^{2}-\tfrac{11\lambda_{k}\rho}{2}\|\Delta\mathbf{z}_{k-1}\|^{3}\right)$
@@ 4.1 Algorithmic scheme
This type of problem is common in the context of optimization and machine learning (Shalev-Shwartz and Ben-David, 2014; Roosta-Khorasani et al., 2014; Roosta-Khorasani and Mahoney, 2019). Thanks to its finite-sum structure, both uniform and nonuniform subsampling schemes satisfy Condition 4.1 with high probability; indeed, we have the stronger theoretical guarantee $\|J(\hat{\mathbf{z}}_{k})-DF(\hat{\mathbf{z}}_{k})\|\leq\tau_{k}$, which implies one of two inequalities in Condition 4.1. As a consequence of Theorem 4.1, our subsampled Newton method achieves the order-optimal iteration complexity of $\Theta(\epsilon^{-2/3})$ for solving the finite-sum convex-concave min-max optimization problems; see Theorem 4.13 for the details.
@@ 4.2 Convergence analysis
We provide our results on the iteration complexity of Algorithm 2 in the following theorem.
@@
$0\leq\tau_{k}\leq\min\{\tau_{0},\tfrac{\rho(1-\kappa_{m})}{4(\kappa_{J}+6\rho)}\|F(\hat{\mathbf{z}}_{k})\|\},\textnormal{ for all }k\geq 0$
@@
Theorem 4.1 shows that Algorithm 2 achieves the same iteration complexity as Algorithm 1 and is thus order-optimal regardless of inexact second-order information and inexact subproblem solving under Conditions 4.1 and 4.2.
In the subsequent analysis, we use the same Lyapunov function: $\mathcal{E}_{t}=\tfrac{1}{2}\|\hat{\mathbf{z}}_{t}-\mathbf{z}_{0}\|^{2}$. The first lemma gives a descent inequality which is analogous to that in Lemma 3.3.
@@
$0<\tau_{k}\leq\min\{\tau_{0},\tfrac{\rho(1-\kappa_{m})}{4(\kappa_{J}+6\rho)}\|F(\hat{\mathbf{z}}_{k})\|\},\textnormal{ for all }k\geq 0$
@@
The first and second terms are bounded using Eq. (4.5) and (4.6). For the third term, we derive from Condition 4.1 that $\|(DF(\hat{\mathbf{z}}_{k-1})-J(\hat{\mathbf{z}}_{k-1}))\Delta\mathbf{z}_{k-1}\|\leq\tau_{k-1}\|\Delta\mathbf{z}_{k-1}\|$. The fourth term is then bounded using the same argument from the proof of Lemma 3.3. Putting these pieces together yields that
$(\mathbf{z}_{k}-\hat{\mathbf{z}}_{k})^{\top}F(\mathbf{z}_{k})$$\leq$ $\|\mathbf{z}_{k}-\hat{\mathbf{z}}_{k}\|\left(\left(\tfrac{\rho}{2}+\kappa_{m}\right)\|\Delta\mathbf{z}_{k-1}\|^{2}+\tau_{k-1}\|\Delta\mathbf{z}_{k-1}\|\right)$
$-6\rho\|\Delta\mathbf{z}_{k-1}\|^{3}+6\rho\|\Delta\mathbf{z}_{k-1}\|^{2}\|\hat{\mathbf{z}}_{k-1}-\hat{\mathbf{z}}_{k}\|.$
@@
Since the descent inequalities in Lemmas 3.3 and 4.3 are the same, Lemma 3.4 and 3.5 also hold true for Algorithm 2. As such, we can apply the same argument used for proving Theorem 3.1 and have

### v5

@@ 1.2 Notation and Organization
[…]
The remainder of the paper is organized as follows. In Section 2, we present the setup of min-max optimization and provide the definitions and optimality criteria. In Section 3, we present an explicit yet conceptual second-order min-max optimization method and prove that it can achieve an iteration complexity of $\Theta(\epsilon^{-2/3})$. In Section 4, we propose a new class of second-order min-max optimization methods with inexact second-order information and inexact subproblem solving, and we provide a crisp characterization of the complexity of solving each subproblem. These proposed inexact methods also inspire a new class of subsampled Newton methods for solving the finite-sum min-max optimization problems. In Section 5, we present the results of experiments on real and synthetic data to demonstrate the practical efficiency of the proposed methods.
@@ 2 Preliminaries
A function $f$ is $\rho$-Hessian Lipschitz if $\|\nabla^{2}f(\mathbf{z})-\nabla^{2}f(\mathbf{z}^{\prime})\|\leq\rho\|\mathbf{z}-\mathbf{z}^{\prime}\|$ for $\forall\mathbf{z},\mathbf{z}^{\prime}$.
@@
[…]
We define the notion of global saddle points for the problem in Eq. (1.1).
@@
[…]
@@
[…]
where $\beta$ satisfies that $\|\hat{\mathbf{z}}-\mathbf{z}^{\star}\|\leq\beta$. Clearly, we have $\mathrm{gap}(\hat{\mathbf{z}},\beta)\geq 0$ since $f(\mathbf{x}^{\star},\hat{\mathbf{y}})\leq f(\hat{\mathbf{x}},\mathbf{y}^{\star})$.
@@
[…]
@@
In our convergence analysis, we define $\mathbf{z}=(\mathbf{x},\mathbf{y})\in\mathbb{R}^{m+n}$ and the operator $F:\mathbb{R}^{m+n}\mapsto\mathbb{R}^{m+n}$:
@@
In the following lemma, we provide the properties of $F$ and its Jacobian $DF$ under Assumptions 2.4 and 2.5. Most of the results in the following lemma are known (Nemirovski, 2004) so we omit the proofs.
@@
Suppose that Assumptions 2.4 and 2.5 hold. Then, we have
@@
    DFDF is $\rho$-Lipschitz continuous, i.e., $\|DF(\mathbf{z})-DF(\mathbf{z}^{\prime})\|\leq\rho\|\mathbf{z}-\mathbf{z}^{\prime}\|$.
@@
$(DF(\mathbf{z})-DF(\mathbf{z}^{\prime}))\textbf{h}=\begin{bmatrix}I_{m}&\\ &-I_{n}\end{bmatrix}(\nabla^{2}f(\mathbf{z})-\nabla^{2}f(\mathbf{z}^{\prime}))\textbf{h}.$ (2.4)
This implies that $\|(DF(\mathbf{z})-DF(\mathbf{z}^{\prime}))\textbf{h}\|=\|(\nabla^{2}f(\mathbf{z})-\nabla^{2}f(\mathbf{z}^{\prime}))\textbf{h}\|$. Thus, we have
$\|DF(\mathbf{z})-DF(\mathbf{z}^{\prime})\|=\sup_{\textbf{h}\neq 0}\left\{\tfrac{\|(DF(\mathbf{z})-DF(\mathbf{z}^{\prime}))\textbf{h}\|}{\|\textbf{h}\|}\right\}=\sup_{\textbf{h}\neq 0}\left\{\tfrac{\|(\nabla^{2}f(\mathbf{z})-\nabla^{2}f(\mathbf{z}^{\prime}))\textbf{h}\|}{\|\textbf{h}\|}\right\}.$
@@ 3 Conceptual Algorithm and Convergence Analysis
As a warm-up, we describe the scheme of Newton-MinMax which is a second-order version of the method in Lin and Jordan (2025) for min-max optimization and yields an iteration complexity of $\Theta(\epsilon^{-2/3})$. We again emphasize that Newton-MinMax is a conceptual algorithmic framework in the sense that it requires exact second-order information and requires the cubic regularized subproblem to be solved exactly.
@@ 3.1 Algorithmic scheme
Our method is a generalization of the first-order extragradient method. Indeed, the $k^{\textnormal{th}}$ iteration consists of two important algorithmic components: (1) we compute $\mathbf{z}_{k+1}=\hat{\mathbf{z}}_{k}+\Delta\mathbf{z}_{k}$ where $\Delta\mathbf{z}_{k}\in\mathbb{R}^{m}\times\mathbb{R}^{n}$ is an exact solution of the nonlinear equation problem given by
$F(\hat{\mathbf{z}}_{k})+DF(\hat{\mathbf{z}}_{k})\Delta\mathbf{z}_{k}+6\rho\|\Delta\mathbf{z}_{k}\|\Delta\mathbf{z}_{k}=\textbf{0};$ (3.1)
(2) we compute $\hat{\mathbf{z}}_{k+1}=\hat{\mathbf{z}}_{k}-\lambda_{k+1}F(\mathbf{z}_{k+1})$.
As suggested by Lin and Jordan (2025), we choose to update $\lambda_{k}$ in an adaptive manner and prove that our method can achieve an iteration complexity of $O(\epsilon^{-2/3})$ under Assumptions 2.4 and 2.5. Such a strategy makes sense; indeed, $\lambda_{k}$ is the step size and would be better to increase as the iterate $\mathbf{z}_{k}$ approaches the set of global saddle points where the value of $\|\Delta\mathbf{z}_{k}\|$ measures the closeness. From a practical viewpoint, Algorithm 1 serves as an alternative to the current pipeline of line-search-based methods – which it simplifies by removing the need for an implicit binary search.
@@ 3.2 Convergence analysis
We provide the iteration complexity of Algorithm 1 as follows,
@@
The key step is to define a Lyapunov function for Algorithm 1: $\mathcal{E}_{t}=\tfrac{1}{2}\|\hat{\mathbf{z}}_{t}-\mathbf{z}_{0}\|^{2}$, which can be used to prove technical results that pertain to the convergence property of Algorithm 1. The first lemma gives us a descent inequality.
@@
$\sum_{k=1}^{t}\lambda_{k}(\mathbf{z}_{k}-\mathbf{z})^{\top}F(\mathbf{z}_{k})\leq\mathcal{E}_{0}-\mathcal{E}_{t}+\underbrace{\sum_{k=1}^{t}\lambda_{k}(\mathbf{z}_{0}-\mathbf{z})^{\top}F(\mathbf{z}_{k})}_{\textbf{I}}+\underbrace{\sum_{k=1}^{t}\lambda_{k}(\mathbf{z}_{k}-\hat{\mathbf{z}}_{k})^{\top}F(\mathbf{z}_{k})-\tfrac{1}{2}\|\hat{\mathbf{z}}_{k}-\hat{\mathbf{z}}_{k-1}\|^{2}}_{\textbf{II}}.$ (3.3)
@@
II $\leq$ $\sum_{k=1}^{t}\left(\tfrac{13\lambda_{k}\rho}{2}\|\Delta\mathbf{z}_{k-1}\|^{2}\|\hat{\mathbf{z}}_{k-1}-\hat{\mathbf{z}}_{k}\|-\tfrac{1}{2}\|\hat{\mathbf{z}}_{k-1}-\hat{\mathbf{z}}_{k}\|^{2}-\tfrac{11\lambda_{k}\rho}{2}\|\Delta\mathbf{z}_{k-1}\|^{3}\right)$
@@ 4.1 Algorithmic scheme
This type of problem is common in optimization and machine learning (Shalev-Shwartz and Ben-David, 2014; Roosta-Khorasani et al., 2014; Roosta-Khorasani and Mahoney, 2019). Thanks to its finite-sum structure, both uniform and nonuniform subsampling schemes satisfy Condition 4.1 with high probability; indeed, we have the stronger theoretical guarantee $\|J(\hat{\mathbf{z}}_{k})-DF(\hat{\mathbf{z}}_{k})\|\leq\tau_{k}$, which implies one of two inequalities in Condition 4.1. As a consequence of Theorem 4.1, our subsampled Newton method can achieve the same iteration complexity of $\Theta(\epsilon^{-2/3})$ for solving the finite-sum convex-concave min-max optimization problems; see Theorem 4.13 for the details.
@@ 4.2 Convergence analysis
We provide the iteration complexity of Algorithm 2 as follows,
@@
$0\leq\tau_{k}\leq\min\{\tau_{0},\tfrac{\rho(1-\kappa_{m})}{4(\kappa_{J}+6\rho)}\|F(\hat{\mathbf{z}}_{k})\|\},\textnormal{ for all }k\geq 0.$
@@
Theorem 4.1 demonstrates that Algorithm 2 achieves the same iteration complexity as Algorithm 1 regardless of inexact second-order information and inexact subproblem solving under Conditions 4.1 and 4.2. Unfortunately, we are unable to claim optimality of Algorithm 2 since existing lower bounds (Adil et al., 2022; Lin and Jordan, 2025) have been established under the assumption of an exact second-order oracle and exact subproblem solutions. When moving to inexact methods, the oracle needs to be redefined. We want to emphasize that there may not be a universal way to define the oracle in this context, particularly when considering matrix operations. For instance, quasi-Newton and subsampled Newton methods are both categorized as inexact Newton methods, yet they differ in how second-order information is approximated, which impacts how their oracles should be defined and, consequently, how lower bounds are characterized. Recent work has begun to address this issue. Specifically, Agafonov et al. (2024a) proposed an accelerated stochastic second-order method that is robust to both gradient and Hessian inexactness, and established a lower bound under a $\delta$-inexact Hessian condition: $\|(H(\mathbf{x})-\nabla^{2}f(\mathbf{x}))(\mathbf{x}^{\prime}-\mathbf{x})\|\leq\delta\|\mathbf{x}^{\prime}-\mathbf{x}\|$ for all $\mathbf{x},\mathbf{x}^{\prime}$. Proceeding a further step, Agafonov et al. (2024b) extended the above results to the VI setting by considering a $\delta$-inexact Jacobian condition: $\|(J(\mathbf{z})-DF(\mathbf{z}))(\mathbf{z}^{\prime}-\mathbf{z})\|\leq\delta\|\mathbf{z}^{\prime}-\mathbf{z}\|$ for all $\mathbf{z},\mathbf{z}^{\prime}$. They proved a lower bound and also provided a thorough comparison with our method (see pages 3-4). Notably, their lower bound is not valid for our inexact method since we bound the Jacobian inexactness with $\tau_{k}>0$ that decreases as $k\rightarrow+\infty$ and bounds the norm of Jacobian from above.
In the subsequent analysis, we use the Lyapunov function: $\mathcal{E}_{t}=\tfrac{1}{2}\|\hat{\mathbf{z}}_{t}-\mathbf{z}_{0}\|^{2}$. The first lemma gives a descent inequality which is analogous to that in Lemma 3.2.
@@
$0<\tau_{k}\leq\min\{\tau_{0},\tfrac{\rho(1-\kappa_{m})}{4(\kappa_{J}+6\rho)}\|F(\hat{\mathbf{z}}_{k})\|\},\textnormal{ for all }k\geq 0.$
@@
The first and second terms are bounded using Eq. (4.5) and (4.6). For the third term, we derive from Condition 4.1 that $\|(DF(\hat{\mathbf{z}}_{k-1})-J(\hat{\mathbf{z}}_{k-1}))\Delta\mathbf{z}_{k-1}\|\leq\tau_{k-1}\|\Delta\mathbf{z}_{k-1}\|$. The fourth term is bounded using the same argument from the proof of Lemma 3.2. Putting these pieces together yields that
$(\mathbf{z}_{k}-\hat{\mathbf{z}}_{k})^{\top}F(\mathbf{z}_{k})\leq\|\mathbf{z}_{k}-\hat{\mathbf{z}}_{k}\|\left((\tfrac{\rho}{2}+\kappa_{m})\|\Delta\mathbf{z}_{k-1}\|^{2}+\tau_{k-1}\|\Delta\mathbf{z}_{k-1}\|\right)-6\rho\|\Delta\mathbf{z}_{k-1}\|^{3}+6\rho\|\Delta\mathbf{z}_{k-1}\|^{2}\|\hat{\mathbf{z}}_{k-1}-\hat{\mathbf{z}}_{k}\|.$ (4.7)
@@
Since the descent inequalities in Lemmas 3.2 and 4.3 are the same, Lemma 3.3 and 3.4 also hold true for Algorithm 2. As such, we can apply the same argument used for proving Theorem 3.1 and have